today the father is 51 years and his son is 9 years . In how many years the father will be four times as old as his son?
Found 2 solutions by ikleyn, greenestamps:
Answer by ikleyn(52788)

.

51 + x = 4*(x+9)


51 + x = 4x + 36

51 - 36 = 4x - x

15      = 3x

x       = 15/3 = 5.


ANSWER.  In 5 years.


Answer by greenestamps(13200)



The difference in their ages is 51-9 = 42 years.

When the father is 4 times as old as his son, their ages will be 4x and x; the difference will be 3x.

3x = 42
x = 42/3 = 14

The father will be 4 times as old as his son when the son is x=14 and the father is 4x=56.

Since the father is now 51, that will be 5 years from now.